OpenStudy (anonymous):

(2y-1) (4y-3) 1) -2y+3 2) 8y^2-10y-3 3) 8y^2-10y+3 4) 8y^2-2y+3 Last problem I need help with right now. I think it is 4 but I am not positive. I believe from what hockeychick23 has told me you combine the like terms and go from there so I am at 8y-4 and I know I have done something wrong I just don't know what it is.

OpenStudy (hockeychick23):

2y*4y= ?

OpenStudy (anonymous):

8y^2

OpenStudy (hockeychick23):

good, now 2y*-3=?

OpenStudy (anonymous):

-6y?

OpenStudy (hockeychick23):

yepp, now -1*4y= ?

OpenStudy (anonymous):

-4y

OpenStudy (hockeychick23):

yepp, now -1*-3= ?

OpenStudy (anonymous):

3

OpenStudy (hockeychick23):

good so now we combine like terms: 8y^2-6y-4y+3= ?

OpenStudy (anonymous):

8y^2-2y+3 ?

OpenStudy (hockeychick23):

not quite, -6y-4y= ?

OpenStudy (♪chibiterasu):

Remember, a negative plus a negative grows.

OpenStudy (anonymous):

10y

OpenStudy (♪chibiterasu):

Just like adding a positive to a positive does the same.

OpenStudy (hockeychick23):

correct! so we have 8y^2-10y+3 as your final answer :)

OpenStudy (anonymous):

YAY Thank you guys so so much

OpenStudy (hockeychick23):

anytime :)

OpenStudy (anonymous):

I got 100%

OpenStudy (hockeychick23):

awesome!